Mr Joe has 3 children Blessing,Elvis and emila given that the second is older than the first one for 3 years and the third is 2times the age of the second calculate their ages if the total sum of their ages is 29
Found 2 solutions by josgarithmetic, ankor@dixie-net.com:
Answer by josgarithmetic(39617)

CHILD     AGE
 B         B
 E        B+3
 L       2(B+3)

SUM        29

Form the simple age sum equation and solve for B.

(In fact, if you look at and think about the expressions carefully, you may solve for B without writing anything on paper.)

Answer by ankor@dixie-net.com(22740)

Mr Joe has 3 children Blessing,Elvis and emila
a, b, c
given that the second is older than the first one for 3 years
b = a+3
or
a = b-3
and the third is 2times the age of the second
c = 2b
calculate their ages if the total sum of their ages is 29
a + b + c = 29
replace a with (b-3), replace c with 2b
(b-3) + b + 2b = 29
combine like terms
4b - 3 = 29
4b = 29 + 3
4b = 32
b = 32/4
b = 8 is Elvis's age
then
a = 8 - 3
a = 5 is Blessing's age
and
c = 2(8)
c - 16 is Emila's age
:
:
Check: 5 + 8 + 16 = 29